OpenStudy (luigi0210):

Find the Derivative:

OpenStudy (luigi0210):

\[y=\frac{3}{(5x^2+\sin2x)^{3/2}}\]

OpenStudy (jdoe0001):

\(\bf \Large y=\frac{3}{(5x^2+\sin2x)^{\frac{3}{2}}} \implies y = 3 \times (5x^2+\sin2x)^{-\frac{3}{2}}\)

OpenStudy (luigi0210):

So I end up with this: \[y'=3*\frac{-3}{2}(5x^3+\sin2x)^{-1/2}*10x+2\cos2x\] Unsimplified, right?

OpenStudy (jdoe0001):

yeah, I'd leave it unsimplified

OpenStudy (luigi0210):

Alright thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

though I left it as \(\bf\Large -\cfrac{9}{2}\left(5x^2+sin(2x)\right)^{-\frac{5}{2}}\)

OpenStudy (anonymous):

use the quotient rule...$$y=3(5x^2+\sin2x)^{-3/2}\\y'=3\times-3/2~(5x^2+\sin2x)^{-5/2}(10x+2\cos2x)=-\frac{9(10x+2\cos2x)}{2(5x^2+\sin2x)^{5/2}}$$

OpenStudy (luigi0210):

Darn, I'm confusing myself with integrals..